Use the Pythagorean theorem to find the missing side.
c=17, b=2​a; find a and b

Answer by richwmiller(17219)

x^2+y^2=z^2
Assuming you mean c to be the hypotenuse
a^2+(2a)^2=17^2
a^2+4a^2=17^2
5a^2=17^2
a^2=17^2/5
a=17/sqrt(5)
a=7.6026311235
b=2a=15.205262247
check
7.6026311235^2+15.205262247^2=17^2
57.8+231.2=289
289=289
ok